Project

General

Profile

NDOS CommissionMtng 20120813

Attending: Denis, Peter, Kurt, Mike Lingren, Zukai, Leon, Fernanda

Installation:
5 fail silicon APD were removed from the detector, 2 were found wet
Leon reports the wet APDs had plugs CB holes
2/3 Dry failed units test out as good on the Caltech QA machine
Black gunk appears to be ash around HV traces
Leon suspects mechanical interference causing the non wet failure
Leon is at FNAL to help in removal and inspection of additional failed APDs
New parylene install ~ 2 weeks out now

Data Check: (Jan was away but now has a script which produces the report below)

Run: 14020
From: 2012-08-03 15:08:21
To: 2012-08-04 15:06:21
Subruns: 24
FEB2-3:0 shutoff in subrun 9

Run: 14021
From: 2012-08-04 15:06:22
To: 2012-08-04 16:02:11
Subruns: 4
There were no shutoffs in this run
The run died in subrun 0

Run: 14022
From: 2012-08-04 20:06:50
To: 2012-08-05 20:04:50
Subruns: 24
FEB3-1:60 shutoff in subrun 0
FEB3-1:3 shutoff in subrun 4

Run: 14023
From: 2012-08-05 20:04:50
To: 2012-08-06 10:05:43
Subruns: 15
There were no shutoffs in this run

Run: 14025
From: 2012-08-06 21:45:46
To: 2012-08-07 08:22:10
Subruns: 11
FEB3-1:3 shutoff in subrun 0
FEB3-1:30 shutoff in subrun 0
FEB3-1:62 shutoff in subrun 0
FEB3-2:59 shutoff in subrun 0

Run: 14028
From: 2012-08-07 22:40:45
To: 2012-08-08 09:05:16
Subruns: 11
FEB1-1:38 shutoff in subrun 1
FEB3-1:43 shutoff in subrun 4
FEB3-1:13 shutoff in subrun 5
FEB2-3:11 shutoff in subrun 6
FEB2-3:32 shutoff in subrun 6

Run: 14030
From: 2012-08-08 14:32:33
To: 2012-08-09 08:55:52
Subruns: 19
FEB3-1:13 shutoff in subrun 0
FEB3-1:36 shutoff in subrun 0
FEB3-1:43 shutoff in subrun 3
FEB2-3:4 shutoff in subrun 12

Run: 14032
From: 2012-08-09 17:02:29
To: 2012-08-10 16:17:15
Subruns: 24
FEB2-1:61 shutoff in subrun 0
FEB3-1:1 shutoff in subrun 0
FEB3-1:6 shutoff in subrun 0
FEB3-1:13 shutoff in subrun 0
FEB3-1:34 shutoff in subrun 0
FEB3-3:32 shutoff in subrun 0
FEB2-1:62 shutoff in subrun 2
FEB2-3:11 shutoff in subrun 2
FEB3-1:43 shutoff in subrun 2
FEB1-1:38 shutoff in subrun 4
FEB3-1:26 shutoff in subrun 4
FEB2-3:13 shutoff in subrun 13
FEB3-3:36 shutoff in subrun 19

Run: 14035
From: 2012-08-10 20:24:51
To: 2012-08-11 20:22:51
Subruns: 22
FEB3-1:6 shutoff in subrun 0
FEB3-1:7 shutoff in subrun 0
FEB3-1:34 shutoff in subrun 0
FEB3-1:48 shutoff in subrun 1
FEB3-1:1 shutoff in subrun 4
FEB3-2:51 shutoff in subrun 9
FEB3-1:26 shutoff in subrun 17

Run: 14036
From: 2012-08-11 20:22:51
To: 2012-08-12 20:20:52
Subruns: 24
FEB3-1:54 shutoff in subrun 21

Run: 14037
From: 2012-08-12 20:20:52
To: 2012-08-13 06:21:01
Subruns: 10
FEB2-3:11 shutoff in subrun 3

Running status:
On Friday high data rate were reported prompting Denis to restart the run
3-1 didn't produce DSO data until a power cycle was performed (rebooting did nothing)
The rates afterward looked about the same
Peter correlated the high rate to new masks loaded at 5 pm on Thurdays (a mass APD recovery set)

DCS- Steve Foulkes requested time for IOC work (This will happen Thursday or Friday)

Zukai - Requested extended DSO scan on behalf on calibration group (he'll work with Andrew and Chris to develop the tools the an exact Run plan)